A gold mining firm sells futures contracts worth 1000 ounces at a price of $700 per ounce for maturity one year from today. If gold futures prices increase to $702 per ounce, what is the cash flow to the producer? O A. $702,000 ○ B. -S2000 C. $700,000 O D. $0 O E. $2000
Answer #1
Sale price = 700
Contract size = 1000
Sale value =700*1000 700000
Gold price at maturity/Buy price = 702
Buy value = 702*1000 702000
Therefore cash flow = 700000-702000 -2000
correct answer is option B. -2000